# Effect of Synthetic Jets Actuator Parameters on Deep Reinforcement Learning-Based Flow Control Performance in a Square Cylinder

### v2

@@ IV CONCLUSIONS
In this study, we investigate the application of DRL-based AFC in mitigating the lift and drag coefficients of a square cylinder while suppressing vortex shedding in the wake field. We design an interaction framework between the agent and a numerical simulation environment, optimizing its control strategy through learning and trial-and-error. The agent observes the state information of the numerical simulation environment and adjusts the mass flow rates of two synthetic jets on the square cylinder based on feedback from reward signals. Through an iterative optimization process, the SAC agent gradually converges towards the optimal flow control strategy, achieving multiple control objectives such as reducing lift and drag coefficients and suppressing vortex shedding. We will evaluate the performance of the SAC agent in controlling synthetic jets on a square cylinder for active flow control, and further analyze the impact of the position and width of the synthetic jet actuators on flow control effectiveness. The main findings of this study can be summarized as follows:
Placing the synthetic jets at both the leading and trailing edge corners significantly reduces the lift and drag coefficients of the square cylinder. However, complete suppression of vortices in the wake field, along with more stable lift and drag coefficients, is only achieved when the synthetic jets are positioned near the trailing edge corners. When the synthetic jets are located at the trailing edge corners, at $Re=100$, the mean drag coefficient decreases by 14.4%, the standard deviation of the drag coefficient reduces by 99%, and the standard deviation of the lift coefficient decreases by 86.1%. Additionally, the vortices in the wake field of the baseline flow are completely suppressed, with no new vortices shedding. At $Re=500$, the mean drag coefficient decreases by 51.4%, the standard deviation of the drag coefficient reduces by 90.5%, and the standard deviation of the lift coefficient decreases by 74.9%. Similarly, the vortices shed in the wake field of the baseline flow are entirely suppressed. Although there is still slight instability in the wake field, no new vortices shed downstream of the square cylinder.
Additionally, we designed four different jet widths with the synthetic jets positioned at the trailing edge corners. The DRL training results indicate that at $Re=100$, with jet widths of $D/20$, $D/25$, $D/30$, and $D/35$, the mean drag coefficient decreases by 13.1%, 14.4%, 13.2%, and 15.2%, respectively. The standard deviation of the drag coefficient reduces by 99% across all widths, while the standard deviation of the lift coefficient decreases by 84.3%, 86.1%, 79.0%, and 68.5%, respectively. This shows that DRL-based AFC effectively reduces both drag and lift coefficients across all jet widths, although narrower jet widths result in more unstable training outcomes, with higher means and standard deviations of the action. At $Re=500$, with jet widths of $D/20$, $D/25$, $D/30$, and $D/35$, the mean drag coefficient decreases by 50.4%, 51.4%, 53.0%, and 48.9%, respectively. The standard deviation of the drag coefficient reduces by 94.8%, 90.5%, 81.8%, and 50.3%, while the standard deviation of the lift coefficient decreases by 84.2%, 74.9%, 75.2%, and 50.3%, respectively. These results demonstrate that both drag and lift coefficients are effectively reduced across all jet widths. However, the reduction rates of the standard deviations for $C_{D}$, $C_{L}$, and the action decrease as the synthetic jets width narrows. When AFC uses narrower synthetic jets, the stability of the controlled flow’s $C_{D}$, $C_{L}$, and action is poorer. This study provides valuable references for the width and position of zero-net mass flux-driven synthetic jets, offering guidance for the practical application of DRL-based active flow control of square cylinders.
@@ ACKNOWLEDGMENTS
The authors would like to express their gratitude to Dr. Jean Rabault (University of Oslo, Oslo, Norway), Dr. Jichao Li (National University of Singapore, Singapore), and Mr. Qiulei Wang (The University of Hong Kong, Hong Kong SAR, China) for making their open-source codes for deep reinforcement learning and numerical simulation available online. These resources have significantly contributed to the research presented in this paper. The relevant code repositories can be accessed at https://github.com/jerabaul29/Cylinder2DFlowControlDRLParallel Rabault and Kuhnle (2019b); Rabault et al. (2019), https://github.com/npuljc/RL_control_Nek5000 Li (2021); Li and Zhang (2022), and https://github.com/venturi123/DRLinFluids Qiulei Wang (2022); Wang et al. (2022).

### v3

@@ IV CONCLUSIONS
This study employs CFD as the training environment for a DRL framework to conduct an AFC investigation on the mass flow rate of synthetic jets on the upper and lower surfaces of a square cylinder, with a focus on examining the influence of the position and width of the synthetic jets on the flow control performance. Through iterative optimization, the DRL agent, specifically a SAC agent, is found to be capable of achieving significant reductions in the lift and drag coefficients of a square cylinder, as well as completely suppress vortex shedding in the wake. The main findings of the study are:
  1. 1.
    DRL-based control is shown to effectively control flow around square cylinders at two different Reynolds numbers of 100 and 500. At $Re$ = 100, synthetic jets located near the trailing edge corners achieve the maximum drag reduction effect, with the mean drag coefficient reduced by 14.4%, the standard deviation of the drag coefficient reduced by 99%, and the standard deviation of the lift coefficient reduced by 86.1%. At $Re$ = 500, synthetic jets located near the leading edge corners achieve the maximum drag reduction effect, with the mean drag coefficient reduced by 51.4%, the standard deviation of the drag coefficient reduced by 90.5%, and the standard deviation of the lift coefficient reduced by 74.9%.
  2. 2.
    In all numerical experiments, the DRL-based flow control strategy enables dynamic adjustments to variations in the physical field. In the beginning, a large jet velocity is often used to rapidly intervene and adjust the flow state. Subsequently, a smaller synthetic jet velocity is employed to sustain the flow control effect and reduce energy consumption. The overall control strategy demonstrates the ability of DRL control in achieving drag reduction while minimizing energy consumption under various flow conditions.
  3. 3.
    The control effectiveness is found to be highly sensitive to the location of synthetic jets Positioning synthetic jets near the trailing corners is advantageous for drag reduction at Re = 100. In contrast, at Re = 500, positioning synthetic jets near the leading corners yields a more pronounced drag reduction effect. Detailed investigations of control mechanisms reveal that optimal control is achieved when the synthetic jets are located at the primary flow separation points. At primary flow separation points, the synthetic jet can most effectively modify the behavior of the boundary layers, directly influencing the location and process of flow separation. By executing blowing or suction actions, the synthetic jets introduce small-scale vortical structures that interfere with the separation bubble to achieve flow control. Without the feeding flow physics into the DRL algorithm, the DRL agent is capable of learning the precise intervention that optimizes the flow control objectives adaptively.
  4. 4.
    The control effectiveness of drag reduction is found to be relatively insensitive to the width of synthetic jets. However, narrower jets can lead to decreased flow stability, particularly at higher Reynolds numbers. A narrower jet concentrates the injected or extracted fluid in a smaller area, resulting in uneven momentum distribution within the boundary layer, which affects the overall flow stability. At $Re=100$, when the jets width decreases from $D/20$ to $D/35$, the standard deviation of the lift coefficient is reduced by 15.8%. When the jet width decreases from $D/20$ to $D/35$ at $Re=500$, the standard deviation of the drag coefficient is reduced by 44.5%, and the standard deviation of the lift coefficient is reduced by 33.9%.
DRL-based AFC offers significant advantages in handling complex nonlinear systems and high-dimensional state spaces, providing powerful tools for flow control. DRL employs a reward mechanism to identify efficient control strategies that significantly reduce lift and drag coefficients and suppress vortex shedding. Investigating the impact of the position and width of synthetic jets on flow control performance can further optimize control strategies, enhance flow stability, and provide valuable insights for practical engineering applications.
@@ ACKNOWLEDGMENTS
The authors would like to express their gratitude to Dr. Jean Rabault (University of Oslo, Oslo, Norway), Dr. Jichao Li (National University of Singapore, Singapore), and Mr. Qiulei Wang (The University of Hong Kong, Hong Kong SAR, China) for making their open-source codes for deep reinforcement learning and numerical simulation available online. These resources have significantly contributed to the research presented in this paper. The relevant code repositories can be accessed at https://github.com/jerabaul29/Cylinder2DFlowControlDRLParallel Rabault and Kuhnle (2019b); rabault2019artificial, https://github.com/npuljc/RL_control_Nek5000 Nek5000; Li and Zhang (2022), and https://github.com/venturi123/DRLinFluids Qiulei Wang (2022); Wang et al. (2022).